**************************************************************** * Linux Network Administrators’ Guide, Version 1.0 * Olaf Kirch, ¡okir@monad.swb.de¿ *************************************************************** * Configuration Options (as set in Config.tex): * Generating PostScript(tm) output * Using PostScript font metrics * Generating DVI output * LaTeX2e(beta) compatibility * LaTeX2e compatibility * Oldtimer TeX installation compatibility * LaTeX 2.09 compatibility * Squeezing output to save a few pages ****************************************************************

[11pt,makeidx]report nftimes epsf epic eepic [11pt]report makeidx times epsfig epic eepic latexsym ifthen lotex private linuxdoc2e [11pt,times,epsf,makeidx,lotex,private,linuxdoc]report [11pt,epsf,makeidx,lotex,private,linuxdoc]report [11pt,epic,eepic,makeidx,lotex,private,linuxdoc]report

[11pt,makeidx,lotex,private,linuxdoc]report The Network Administrators’ Guide Olaf Kirch 1992-1994 [] This is the last pre-release before the

new Networking Guide goes public. I consider the technical chapters mostly complete. I have tried to catch up with the current development, but in some section, there have been minor changes or additions to the software since I wrote them. I don’t think I will update them unless you tell me I’ve missed something big.

1em If you have any comments, suggestions, or complaints, please mail me at okir@monad.swb.de. – Olaf 5mm

secnumdepth2 tocdepth2 tex2html_{w}rapdocumentemptysecnumdepth2tocdepth2 minipage0.9 Legal Notice

UNIX is a trademark of Univel. is not a trademark, and has no connection to UNIX or Univel. * Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Olaf Kirch Kattreinstr. 38, 64295 Darmstadt, Germany okir@monad.swb.de

.2in“The Network Administrators’ Guide” may be reproduced and distributed in whole or in part, subject to the following conditions:

enumerate enumi-1 T he copyright notice above and this permission notice must be preserved complete on all complete or

partial copies. A ny translation or derivative work of “The Network Administrators’ Guide” must be approved by the

author in writing before distribution. I f you distribute “The Network Administrators’ Guide” in part, instructions for obtaining the complete

version of “The Network Administrators’ Guide” must be included, and a means for obtaining a complete version provided.

S mall portions may be reproduced as illustrations for reviews or quotes in other works without this

permission notice if proper citation is given. I f you print and distribute “The Network Administrators’ Guide”, you may not refer to it as the “Official

Printed Version”. T he GNU General Public License referenced below may be reproduced under the conditions given within

it. S everal sections of this document are held under separate copyright. When these sections are covered by a

different copyright, the seperate copyright is noted. If you distribute “The Network Administrators’ Guide” in part, and that part is, in whole, covered under a seperate, noted copyright, the conditions of that copyright apply.